OpenStudy (mathmath333):

7 different objects must be divided among 3 people . In how many ways can this be done if at least one of them gets exactly one object ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{7 different objects must be divided among 3 people .}\hspace{.33em}\\~\\ & \normalsize \text{In how many ways can this be done if at least one}\hspace{.33em}\\~\\ & \normalsize \text{of them gets exactly one object ?} \hspace{.33em}\\~\\ & a.)\ 2484 \hspace{.33em}\\~\\ & b.)\ 1218 \hspace{.33em}\\~\\ & c.)\ 729 \hspace{.33em}\\~\\ & d.)\ \normalsize \text{none of these} \end{align}}\)

OpenStudy (anonymous):

i know hw to do it but u do know the detail solution for this comes if u google u want me to give the link and if u have question i ll answer :)

OpenStudy (arindameducationusc):

Combination question?

OpenStudy (mathmath333):

well the fact is that i didnt understand the quora answer

imqwerty (imqwerty):

nd what about the gmat solution :)

OpenStudy (mathmath333):

also the gmat :)

OpenStudy (arindameducationusc):

@phi

OpenStudy (mathmath333):

isnt he offline

OpenStudy (arindameducationusc):

yes, but if he comes online, he will definitely see, Hez good!

OpenStudy (arindameducationusc):

I am trying....

imqwerty (imqwerty):

ok 1st can u find the number of ways in which u can distribute the gifts :)consider all cases

OpenStudy (mathmath333):

7P3 ?

imqwerty (imqwerty):

did u count the possibility in which they don't even get a single gift

OpenStudy (mathmath333):

how to do that

imqwerty (imqwerty):

m also thinkin that

OpenStudy (arindameducationusc):

:)

imqwerty (imqwerty):

(:

OpenStudy (mathmate):

I have the following calculation, but have succeeded in making a combinatorial verification yet. Ways to distribute 7 different objects to 3 different bins = 3^7 (order is not important in each bin) ways to choose 2 bins out of 3 = 3 Ways to distribute 7 different objects to 2 different bins = 2^7 ways to choose 1 bin out of 3 = 3 ways to distribute 7 different objects to 1 bin = 1^7 So total number of ways to distribute 7 different objects to 3 different bins, with at least one object in each bin = 3^7-3*2^7-3*1^7 = 2187-3(128)-3*(1) =1800

OpenStudy (mathmate):

*have not succeeded

OpenStudy (mathmate):

Oh, I did not read correct, it says at least one of them gets \(exactly\) one object. I will do a variation of the above and come back.

Parth (parthkohli):

Hey, I guess I know this. But I'm not sure if we're allowed to give zero objects to any of them. It depends on the case.

OpenStudy (mathmath333):

yes u r allowed to give 0 objects i think.

Parth (parthkohli):

First, we fix the person who gets one gift. That makes us fulfil the initial condition. There are \(3\) ways to choose the person who gets one gift. Now there are six gifts left as we've given one gift to one person.\[x+y = 6\]has solutions \((0,6), (1,5),(2,4),(3,3), (4,2),(5,1),(6,0) \).

OpenStudy (mathmath333):

ok

Parth (parthkohli):

You know what, there's another way to do this.

OpenStudy (mathmath333):

well i just need an understandable way lol

Parth (parthkohli):

OK, I'd done it but the net got disconnected.

Parth (parthkohli):

First consider the case where two people get one object each. First we have to choose two people, in \(\binom{3}{2}\) ways. Then we have to choose 2 objects out of 7 and permute them among these two guys. That can be done in 7P2 ways.

OpenStudy (mathmath333):

3*(7*(2^6-6))

Parth (parthkohli):

Then consider the case where only one guy gets one object. First choose the guy who gets that one object. \(3\) ways. Then the other two guys have (0,6) (2,4) (3,3) (4,2) (6,0)

Parth (parthkohli):

For the first and the last case, we can count it in this way: \[2 \times \binom{7}{6}\]For the second and the second-to-last case, here's how:\[2 \times \binom{7}{4}\times \binom{3}{2}\]For the third case,\[2\times \binom{7}{3}\times \binom{4}{3}\]

Parth (parthkohli):

Heyyyyy never mind all that, you know

Parth (parthkohli):

OK, unfortunately that's it.

Parth (parthkohli):

Have we double-counted?

Parth (parthkohli):

@ganeshie8 our saviour

Parth (parthkohli):

Case 1: Exactly two persons get exactly one object.\[\binom{3}{2}\cdot \binom{7}{2}\cdot 2 \]Case 2: Exactly one person gets exactly one object. Number of ways to choose that person is 3. Now the other two are selected.\[0 ~\& ~6: 2! \cdot \binom{7}{6} \]\[2 ~ \& ~ 4: 2! \cdot \binom{7}{4}\cdot \binom{3}{2}\]\[3 ~\& ~3:\binom{7}{3}\cdot \binom{4}{3 } \]

OpenStudy (mathmath333):

answer given =1218

Parth (parthkohli):

Still not correct.

Parth (parthkohli):

ooo got it

ganeshie8 (ganeshie8):

1) Give "one" person "one" gift and forget him. This can be done in \(3\times 7\) ways. (3 ways to choose 1 person, 7 ways to choose 1 gift) 2) Then the task is to distribute "6" gifts to "2" people such that nobody gets exactly "1" gift : \(2^6-6\) Multiply them for the final answer.

ganeshie8 (ganeshie8):

At step2, above, the expression \(2^6\) comes from the fact that each gift has \(2\) possible states

OpenStudy (mathmath333):

that looks simple now

Parth (parthkohli):

Much sleeker...

Parth (parthkohli):

Haha, looks like mathmath had posted that expression earlier. :P

ganeshie8 (ganeshie8):

Actually it was I who posted it from mathmath's computer..

Parth (parthkohli):

...what

OpenStudy (mathmath333):

He is a hacker

Parth (parthkohli):

oh ew i missed the sarcasm there >_<

ganeshie8 (ganeshie8):

lol we were doing teamviewer earlier but my headset won't cooperate suddenly... so..

Parth (parthkohli):

Why did you subtract six?

OpenStudy (mathmath333):

to substract the possibltties where a person gets exactly 1 gift

ganeshie8 (ganeshie8):

At step2, we have 6 gifts and two people. we don't want to give exactly "1" gift to anyone. Actually, we need subtract \(2*6\), I think. Need to think more...

Parth (parthkohli):

Yeah, I mean you can give exactly two gifts to two people, as long as the last guy gets five.

Parth (parthkohli):

what I mean is you can give exactly one gift to two people

ganeshie8 (ganeshie8):

Yep, so the answer is wrong. It should be : \(3*7(2^6-2*6) = 1092\)

ganeshie8 (ganeshie8):

because, at step2, we have 6 gifts and 2 people each person can get "1" gift in 6 ways, so 2*6 total invalid states

Parth (parthkohli):

What am I even saying ugh

ganeshie8 (ganeshie8):

No, keep going..

Parth (parthkohli):

No, there are six invalid states only!

Parth (parthkohli):

But there are two people, so you're right.

Parth (parthkohli):

But why exactly is (1, 1, 5) being counted as invalid?

ganeshie8 (ganeshie8):

Because the question says so : "exactly one person gets one gift"

Parth (parthkohli):

no, it says at least one person gets exactly one gift

Parth (parthkohli):

which is why we need cases

ganeshie8 (ganeshie8):

Oops! I read the problem wrong, wait

Parth (parthkohli):

Hey, I found this problem's solution. http://prntscr.com/8k0bq7

Parth (parthkohli):

The solution is pretty much how I did it. The only difference was that this solution first distributed one gift to the guy who had one gift, then distributed it to another guy. What I did was that I distributed that one gift in the end.

Parth (parthkohli):

http://prntscr.com/8k0czp

ganeshie8 (ganeshie8):

that is nice

Parth (parthkohli):

\[\binom{3}{2}\cdot \binom{7}{2}\cdot 2\]\[+ ~3 \left(2!\cdot \binom{7}{6} + 2!\cdot \binom{7}{4}\cdot \binom{3}{2} + \binom{7}{3}\cdot \binom{4}{3}\right)\]

ganeshie8 (ganeshie8):

Hmm, wonder how is that related to \(3*7(2^6-6)\)

Parth (parthkohli):

oh that's actually right... I evaluated it wrongly the last time

Parth (parthkohli):

thanks wolfram

Parth (parthkohli):

i was getting 1638 the last time because I forgot there was no 2! for the very last term -_-

Parth (parthkohli):

what were you guys doing on teamviewer anyway